# ShoppingComp: Are LLMs Really Ready for Your Shopping Cart?

### v1

@@ Abstract
We present ShoppingComp, a challenging real-world benchmark for rigorously evaluating LLM-powered shopping agents on three core capabilities: precise product retrieval, expert-level report generation, and safety critical decision making. Unlike prior e-commerce benchmarks, ShoppingComp introduces highly complex tasks under the principle of guaranteeing real products and ensuring easy verifiability, adding a novel evaluation dimension for identifying product safety hazards alongside recommendation accuracy and report quality. The benchmark comprises 120 tasks and 1,026 scenarios, curated by 35 experts to reflect authentic shopping needs. Results reveal stark limitations of current LLMs: even state-of-the-art models achieve low performance (e.g., 11.22% for GPT-5, 3.92% for Gemini-2.5-Flash). These findings highlight a substantial gap between research benchmarks and real-world deployment, where LLMs make critical errors such as failure to identify unsafe product usage or falling for promotional misinformation, leading to harmful recommendations. ShoppingComp fills the gap and thus establishes a new standard for advancing reliable and practical agents in e-commerce. Our code and dataset are available at https://github.com/ByteDance-BandAI/ShoppingComp
Figure 1: Leaderboard comparison on the four evaluation dimensions of ShoppingComp. Top-left: Product retrieval (AnswerMatch-F1). Top-right: Scenario Coverage-F1 for report comprehensiveness. Bottom-left: Report Rationale Validity. Bottom-right: Safety Rubric Pass Rate.
@@ 1 Introduction
The rise of large language models (LLMs) has sparked increasing interest in deploying them as intelligent shopping assistants, capable of retrieving products, generating recommendations, and guiding consumer decisions. Now, with OpenAI’s newly released ChatGPT Shopping Research OpenAI (2025c) feature in ChatGPT, the promise of such assistants is becoming more concrete: users can ask vague, purpose-driven questions (e.g., “find a quiet cordless vacuum for a small apartment”) and receive structured, comparative buyer’s guides synthesized from live web data. However, despite promising results on academic benchmarks, there remains a striking gap between benchmark performance and real-world deployment. Figure. 2 user instance shows consumers typically express needs as multi-constraint problems, where constraints arise from real usage scenarios and everyday contexts. Also, recommending unsafe products or falling prey to promotional misinformation can directly harm users and undermine confidence in AI systems. To build shopping agents that are not only effective but also reliable, evaluation frameworks must capture both the complexity of authentic consumer needs and the safety of consumer decisions.
Existing e-commerce benchmarks only partially capture these requirements (Tab. 1). WebShop (Yao et al., 2022), Mind2Web (Deng et al., 2023) evaluates task-driven browsing in simulated websites, and OPeRA (Wang et al., 2025b) models user–agent interactions. While valuable, these datasets rely on closed-world assumptions, static product sets, noise-free contexts, and limited environmental variability—thus failing to test models under realistic conditions such as changing product availability, misleading marketing content, and vague, goal oriented consumer intents that require open ended reasoning. Beyond shopping, rubric-based evaluation has been explored in other domains; for instance, HealthBench (Arora et al., 2025) underscores the need for domain specific rigor in health-related tasks. Building on these efforts, our work extends rubric-based evaluation to the shopping domain, with an explicit focus on consumer safety and reliability.
@@
Table 1: Comparison with prior shopping/web-agent benchmarks. ShoppingComp uniquely combines real-world products, rubric-driven report evaluation, and safety-critical evaluation under an end-to-end open-world setup. Here, ◗ denotes partially satisfied.
Dimension WebShop Mind2Web OPeRA ShoppingComp (Ours)
Real-world products (live market)
Open-world web search (no fixed products)
Report evaluation (rubric-driven)
@@
Verifiability (URL/image evidence)
Persona
We introduce ShoppingComp, a benchmark built on real, verifiable products and high-complexity tasks grounded in authentic consumer needs. To illustrate its design, Figure. 2 shows representative examples: user-authored, expert-authored, and safety-critical cases. Each task is paired with detailed rubrics, ground-truth products, and verifiable evidence, enabling transparent and reproducible evaluation. ShoppingComp evaluates agents across three complementary tasks: (1) Browse Products. Assesses whether models can retrieve real, commercially available products that meet complex user needs under realistic search noise; (2) Expert-level Report Generation. Evaluates the ability to produce accurate, rubric aligned product reports with verifiable reasoning; (3) Safety-Critical Decision Making. Tests the ability to recognize and avoid product related risks.
Figure 1 summarizes model performance across these tasks and corresponding metrics: AnswerMatchF1 for retrieval, Scenario Coverage and Rationale Validity for report, and Safety Rubric Pass Rate for safety. All models remain far from achieving human-level reliability, with retrieval as the primary bottleneck, reasoning shows moderate gains, and safety evaluation underscores that preventing user harm in real-world deployment remains the most critical unresolved challenge.
  • Realistic and challenging benchmark: We present ShoppingComp, comprising 120 tasks and 1,026 scenarios curated by 35 experts with over 1,000 hours of effort. All tasks employ real search tools under temporal constraints, rule-based validation, and rubric-based grading to ensure realism, difficulty, and verifiability. Alongside the benchmark, we release a dedicated verifier test set to measure the accuracy of LLM-as-a-Judge grading, enabling future verifier model development.
  • Holistic evaluation with safety focus: ShoppingComp jointly evaluates product retrieval, expert-level report generation, and safety-critical decision making, introducing safety-related trap questions and rubric-based checks to assess safety performance.
  • Novel rubric-based report assessment: Beyond product selection, our rubric also evaluates report comprehensiveness, rationale validity, and the inclusion of safety warnings, providing a more faithful measure of real-world reliability.
  • Empirical insights: Experiments further reveal three key issues: weak reasoning over authentic consumer needs, fragile robustness in safety-critical contexts, and the persistent gap between machine and human performance.
@@ 2 Related Work
Web Agent Benchmarks. Early benchmarks such as GAIA (Mialon et al., 2023) target general-purpose assistants, while BrowseComp (Wei et al., 2025) focuses on evaluating browsing skills.
Shopping Agent Benchmarks. E-commerce has recently emerged as a practical application domain. WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023) and Mind2Web (Deng et al., 2023) examine how agents follow natural language instructions to find products, but they rely on simulated environments and primarily assess attribute matching. Shopping-specific datasets such as Product Comparison Corpus (Vedula et al., 2023), ShoppingMMLU (Jin et al., 2024), eCeLLM (Peng et al., 2024), ShoppingBench (Wang et al., 2025a) and OPeRA (Wang et al., 2025b) extend evaluation to multi-task reasoning and user behavior simulation, yet still lack end-to-end testing under realistic and safety-critical conditions.
Rubric-Based Evaluation. Benchmarks like HealthBench (Arora et al., 2025) highlight the need for domain-specific rigor, and recent works (Hashemi et al., 2024; Fan et al., 2025; D’Souza et al., 2025) propose rubric-driven evaluation for fine-grained, interpretable assessment. Our work extends these efforts to the shopping domain by integrating real product search with rubric-based evaluation and explicit safety testing.
@@ 3.1 Task Definition
To address these evaluation gaps, we present the design and construction of ShoppingComp. We design three tasks to cover the full spectrum of shopping agents: Browse Products, Expert-level Report Generation and Safety-Critical Decision Making.
Browse Products. Inspired by BrowseComp, this task evaluates whether models can accurately retrieve real, commercially available products from a noisy and vast search space. It is easy to verify but hard to solve: questions reflect complex real-life needs, where simple attribute matching fails and brute-force search is infeasible. The task therefore stresses efficient strategies and advanced reasoning, highlighting core browsing capabilities.
Expert-level Report Generation. Unlike traditional search, AI shopping agents are expected to produce structured reports explaining product choices. This task assesses reports against expert-defined rubrics, focusing on comprehensiveness, accuracy, and justification. The novelty lies in evaluating not only what products are recommended but also why, making trustworthiness and reasoning quality central to performance.
Safety-Critical Decision Making. A unique contribution of ShoppingComp is the inclusion of safety traps, where experts embed potential hazards into queries. Models are judged on whether they recognize risks and provide appropriate warnings or safe alternatives. This task introduces a critical dimension absent from prior benchmarks, ensuring agents are tested on safety awareness alongside retrieval and reasoning.
@@ 3.2 Data Collection
Step1: Question Collection
@@
Type 1 – Synthesized User Questions. We derive user needs directly from real, currently sold products. Starting with items that feature rich specifications and high decision complexity, we extract key attributes such as capacity, power, and compatibility, and use them to construct multi-constraint shopping scenarios. LLMs generate natural-language queries expressing the underlying intent behind these attributes, which are then verified by human annotators for realism and coherence in the subsequent step.
Type 2 – Expert Questions. Experts distilled consumer dialogues into structured rubrics that encode explicit constraints (e.g., size, standards) and implicit expectations (e.g., durability, usability), ensuring scenarios remain authentic and verifiable.
Type 3 – Safety Questions. A novel contribution of ShoppingComp is the design of safety traps. Experts derived risk-prone scenarios (e.g., unsafe home appliance installation, skin irritation caused by improper use of skincare products), then formalized rubrics requiring compliance with safety codes and hazard awareness. This adds a unique evaluation axis of consumer protection and trust.
@@
Step2: Rubric Generation For each question, both experts and LLMs generate detailed rubrics specifying requirements and standards. Human annotators perform a thorough correctness check for each rubric and consolidate the validated results into the final rubric set. Rubrics act as an intermediate reasoning layer that helps decompose complex e-commerce problems into smaller, scenario-level subproblems, and concrete combinations of product attributes, which greatly reduces task ambiguity and difficulty. For instance, the demand “a rice cooker suitable for a family of three to four” can be logically grounded to a capacity requirement of around 3 L.
@@
Step 3: Multi-source Candidate Product Construction. Candidate products are collected from multiple sources, including web agent retrieval, expert-curated lists and similarity-based retrieval using product embeddings. This stage forms a diverse pool that contains both relevant and misleading candidates. Annotators then link each product to verifiable evidence such as official specifications, trusted reviews or product images. For example, in task on choosing a gaming mouse (Figure. 3), the system retrieved 34 candidate products, but only 3 were verified to fully meet all rubric requirements.
@@
Step 4: Ground-truth Verification and Instance Filtering Human annotators first assess whether candidate products match the rubrics and provide supporting evidence. We then remove trivial or redundant questions through three filters: (1) questions with excessive valid products (#products >> 10), (2) easy questions correctly solved by most evaluated agents, and (3) semantically clustered duplicates identified through embedding-based similarity. The remaining instances undergo expert cross-review to ensure reliability and fairness. This layered filtering process guarantees that final benchmark tasks are challenging, diverse, and rigorously validated.
@@
Shopping experts cohort, selection and input. We involved two teams: a panel of 35 vetted domain experts (1,000 person-hours) and a 15-member annotation team (3,000 person-hours). Experts, recruited and vetted for professional credibility, curated queries across domains, with chairs mediating disagreements. Annotators, though non-specialists, verified correctness, added valid answers, and gathered supporting evidence under expert guidelines. This dual process balanced domain expertise with large-scale verification.
3.3 LLM-as-a-Judge Verification
To reduce manual cost and improve consistency, we introduced two LLM-as-a-Judge verifiers, built on Gemini-2.5-Pro with Google search. Their reliability was validated against dedicated human-annotated test sets, showing strong alignment. More ablation study is shown in Appx. D.
Product Verifier. This verifier checks whether a product satisfies scenario-specific requirements and aggregates judgments at the question level. It reached 81% agreement at the scenario level and 84% at the question level, with discrepancies partly due to human annotation noise and partly to retrieval or reasoning errors.
Report Verifier. This verifier evaluates report quality against rubrics and evidence. It achieved 75.6% agreement with humans—lower than the Product Verifier due to stricter criteria requiring both correct rubric use and reasoning. This highlights the inherent difficulty and subjectivity of report evaluation.
Overall, this dual-verifier design introduces a novel, scalable way to automate benchmark validation while preserving human-level rigor.
3.4 Human Performance And Dataset Diversity
The time distribution for experts and annotators is shown in Figure. 5. We collect response times from both experts and annotators through an online examination. To ensure fair comparison and reduce the impact of anomalous durations, we cap the recorded time: responses longer than one hour for experts and two hours for annotators are excluded from the final analysis. The long completion time underscores the intrinsic difficulty of ShoppingComp tasks, even for humans, establishing a rigorous upper bound for model evaluation.
@@ 3.4 Human Performance And Dataset Diversity
ShoppingComp consists of 120 instances covering 1,026 real-world scenarios, including 54 synthesized user tasks, 40 expert-generated tasks, and 26 safety critical tasks. The benchmark follows the Amazon taxonomy and includes ten categories (see Figure. 5).

### v2

@@ Abstract
We present ShoppingComp, a challenging real-world benchmark for comprehensively evaluating LLM-powered shopping agents on three core capabilities: precise product retrieval, expert-level report generation, and safety critical decision making. Unlike prior e-commerce benchmarks, ShoppingComp introduces difficult product discovery queries with many constraints, while guaranteeing open-world products and enabling easy verification of agent outputs. The benchmark comprises 145 instances and 558 scenarios, curated by 35 experts to reflect authentic shopping needs. Results reveal stark limitations of current LLMs: even state-of-the-art models achieve low performance (e.g., 17.76% for GPT-5.2, 15.82% for Gemini-3-Pro). Error analysis reflects limitations in core agent competencies, including information grounding in open-world environments, reliable verification of multi-constraint requirements, consistent reasoning over noisy and conflicting evidence, and risk-aware decision making. By exposing these capability gaps, ShoppingComp characterizes the trust threshold that AI systems must cross before they can be proactively trusted for reliable real-world decision making. Our code and dataset are available at https://github.com/ByteDance-BandAI/ShoppingComp
Figure 1: Leaderboard comparison on the four evaluation dimensions of ShoppingComp. Top-left: Product Accuracy (AnswerMatch-F1). Top-right: Rubric-based Score of Products (SoP). Bottom-left: Report Rationale Validity. Bottom-right: Safety Rubric Pass Rate.
@@ 1 Introduction
Large Language Models (LLMs) are evolving into autonomous agents for real-world tasks, such as online shopping. Shopping presents a difficult decision making problem, as product information is scattered across diverse web sources and user requirements often involve multiple interacting constraints. With OpenAI’s newly released ChatGPT Shopping Research (OpenAI, 2025d), such assistants are now deployed in real-world products. Users can ask vague, purpose driven questions (e.g., “find a quiet cordless vacuum for a small apartment”) and receive recommendations of real, purchasable products accompanied by structured comparative reports. As shopping assistants evolve, they require more comprehensive evaluation. An ideal benchmark should account for open-world web environments, real product retrieval, and comparative and safety report generation.
As summarized in Table 1, existing shopping benchmarks predominantly adopt closed-world product assumptions. For example, ShoppingBench (Wang et al., 2025a) builds a sandbox with over 2.5 million products, yet still operates under a bounded catalog, abstracting away the noisy, heterogeneous, and often conflicting information commonly encountered in open-world online marketplaces. In contrast, real-world user requests are typically expressed as goal-oriented intents rather than explicit attribute constraints, such as queries like “suitable for elderly use,” which require domain knowledge and open-web search to infer latent product requirements (e.g., weight limits, power ratings, or safety certifications). Such implicit multi-constraint reasoning is not explicitly evaluated in existing shopping benchmarks, such as WebShop (Yao et al., 2022). Moreover, practical shopping assistants must retrieve multiple products that satisfy user needs and provide clear justifications explaining why each option is suitable, in order to support informed decision making, which is partially considered by ProductComparisonCorpus (Vedula et al., 2023). Finally, shopping decisions often involve safety-critical considerations, as incorrect recommendations may cause severe real-world consequences, e.g. failing to recognize that metal containers must not be used in microwave mode, but such failure modes are rarely assessed. ShoppingComp addresses these gaps through end-to-end open world evaluation with rubric-based assessment of reasoning, justification, and safety awareness.
@@
Table 1: Comparison with prior shopping benchmarks.
Evaluation Dimension WebShop ProductComparisonCorpus ShoppingBench ShoppingComp
Open-world products
Open-world web search
Multi-constraint reasoning
Multi-product justification reports
@@
We introduce ShoppingComp, a benchmark built on open-world products and high-complexity shopping tasks. Figure 2 illustrates its design with representative examples, including synthesized user instance, expert instance, and safety instance. Each task is paired with detailed rubrics, ground-truth products, and verifiable evidence to support a reliable evaluation.
ShoppingComp evaluates shopping assistants across three tasks: (1) Browse Products, which assesses whether models can retrieve real products that satisfy complex user needs under search noise; (2) Expert-level Report Generation, which evaluates the ability to rubric-aligned product recommendation reports with verifiable reasoning; and (3) Safety-Critical Decision Making, which tests whether models can recognize and avoid potential product related risks.
Figure 1 summarizes model performance across these tasks and their corresponding metrics, including AnswerMatch F1 and SoP for product accuracy, Rationale Validity for report quality, and Safety Rubric Pass Rate for safety evaluation. All evaluated models remain far from achieving human-level reliability. Our contributions are as follows:
  • A Realistic Expert-Curated Benchmark: We introduce ShoppingComp, comprising 145 instances and 558 scenarios curated by 35 experts with over 1,000 hours of effort. Each task involves complex multi-constraint shopping queries, paired with human-verified product lists, annotated rubrics and verifiable evidence.
  • Unified Evaluation Framework: We introduce a unified evaluation framework that combines AnswerMatch with fine-grained rubrics. The framework supports a comprehensive assessment of LLM performance across four key dimensions: product retrieval accuracy, information acquisition effectiveness, reasoning faithfulness, and safety-critical awareness.
  • Empirical Insights: Our experimental study reveals a significant gap between state-of-the-art LLMs and human experts in open-world shopping. Our analysis uncovers fundamental bottlenecks in large-scale retrieval and information reliability handling, and highlights implicit requirement inference as a dominant and previously underexplored failure mode that propagates errors across retrieval, reasoning, and safety dimensions.
@@ 2 Related Work
Web Agent and Shopping Benchmarks. General-purpose web benchmarks such as WebArena (Zhou et al., 2023) and Mind2Web (Deng et al., 2023) established the foundation for general web navigation, incorporating e-commerce as a key sub-domain. GAIA (Mialon et al., 2023), BrowseComp (Wei et al., 2025) and HLE (Phan et al., 2025) assess high-level tool use and multi-hop information reasoning. Although they include shopping-related tasks, their focus remains on UI interactions and generic information seeking, rather than the core decision-making challenges of real-world e-commerce.
Initial efforts like WebShop (Yao et al., 2022) evaluates instruction following in simulated environments, while ShoppingMMLU (Jin et al., 2024), eCeLLM (Peng et al., 2024) and OPeRA (Wang et al., 2025b) introduce multi-task reasoning and user behavior modeling. Large-scale environments such as ShoppingBench (Wang et al., 2025a) ground evaluation in real product catalogs, but remain restricted to closed-world settings.
Rubric-Based Evaluation. Recent work has adopted rubric-based evaluation to enable fine-grained and interpretable assessment across a wide range of domains, including healthcare (HealthBench (Arora et al., 2025)), scientific reasoning (YESciEval (D’Souza et al., 2025)), and finance and law (PRbench (Akyürek et al., 2025)). This paradigm has also been applied to general question answering and information-seeking tasks, such as ResearchRubrics (Sharma et al., 2025), SedArEval (Fan et al., 2025), and LLM-rubric (Hashemi et al., 2024). Our work extends rubric-based evaluation to the e-commerce domain.
@@ 3.1 Task Definition
We design three tasks to cover the full spectrum of shopping agents.
Browse Products. This task evaluates whether models can retrieve products that satisfy user constraints. For each query, we provide a human-verified product list, enabling straightforward measurement of product-level precision and recall through Answer Match. In addition, fine-grained rubrics are employed to evaluate constraint-level requirement satisfaction. The task is challenging due to complex real-world queries that involve multiple constraints, where simple attribute filtering is insufficient and exhaustive search is infeasible in open world products.
Expert-level Report Generation. Beyond retrieving products, shopping agents are expected to generate structured reports that explain and justify their recommendations. This task evaluates reports using rubrics, focusing on factual accuracy, coverage of user requirements, and the quality of supporting rationales. By assessing not only which products are recommended but also why, this task places reasoning transparency and decision faithfulness at the center of evaluation.
Safety-Critical Decision Making. A distinctive aspect of ShoppingComp is the explicit inclusion of safety-critical scenarios. Experts embed potential hazards and misuse risks into shopping queries, requiring models to identify unsafe conditions and respond with appropriate warnings or safer alternatives. Safety performance is evaluated using a dedicated trap rubric, which explicitly assesses whether models recognize safety risks and take appropriate precautionary actions.
Across all three tasks, each instance is anchored to a user shopping question and includes expert-annotated products, structured rubrics specifying fine-grained requirements and safety conditions, as well as verifiable supporting evidence.
@@ 3.2 Data Collection
Step1: Question Collection. We collect shopping questions from three complementary sources to reflect diverse and realistic consumer needs.
@@
Type 1 – Synthesized User Questions. We begin by selecting real-world products with rich specifications and high decision complexity. Key attributes such as capacity, power, size, compatibility, and supported accessories are extracted to form structured requirement sets. Based on these attributes, LLM-based web agents generate natural-language shopping questions that express the underlying purchase intent behind the constraints. These synthesized questions are subsequently reviewed by human annotators to ensure realism, coherence, and consistency with real consumer expressions.
Type 2 – Expert Questions. To increase data diversity and ensure realism, domain experts directly author shopping questions based on their experience with real consumer needs and purchasing scenarios. These expert-written questions cover a wide range of product categories, usage contexts, and constraint combinations that are difficult to synthesize automatically. They include both explicit constraints (e.g., size limits, technical standards, budget ranges) and implicit requirements (e.g., durability, usability, long-term maintenance), providing complementary coverage to synthesized questions. All expert questions are subsequently passed to Step 2 for rubric generation and verification.
Type 3 – Safety Questions.
To incorporate safety critical decision making, experts curate risk prone scenarios grounded in real usage contexts and domain knowledge (e.g., unsafe home appliance installation, skin irritation caused by improper use of skincare products). These scenarios are formulated into shopping questions that require explicit safety awareness, and are later paired with dedicated safety rubrics during rubric generation and verification.
Across all three types, questions are unified into a shared pool and serve as the input to Step 2 for rubric generation.
@@
Step2: Rubric Generation. For each question, both experts and LLMs generate detailed rubrics specifying requirements and standards. Human annotators perform a thorough correctness check for each rubric and consolidate the validated results into the final rubric set. Rubrics act as an intermediate reasoning layer that helps decompose complex e-commerce problems into smaller, scenario-level subproblems, and concrete combinations of product attributes, which greatly reduces task ambiguity and difficulty. For instance, the demand “a rice cooker suitable for a family of three to four” can be logically grounded to a capacity requirement of around 3 L.
@@
Step 3: Multi-source Candidate Product Construction. Candidate products are collected from multiple sources, including web agent retrieval, expert-curated lists and similarity-based retrieval using product embeddings. This stage forms a diverse pool that contains both relevant and misleading candidates. Annotators then link each product to verifiable evidence such as official specifications, trusted reviews or product images. For example, in task on choosing a gaming mouse (Figure 3), the system retrieved 34 candidate products, but only 3 were verified to fully meet all rubric requirements.
@@
Step 4: Ground-truth Verification and Instance Filtering. Human annotators first assess whether candidate products match the rubrics and provide supporting evidence. We then remove trivial or redundant questions through three filters: (1) questions with excessive valid products (#products >> 10), (2) easy questions correctly solved by most evaluated agents, and (3) semantically clustered duplicates identified through embedding-based similarity. The remaining instances undergo expert cross-review to ensure reliability and fairness. This multi-stage filtering process guarantees that final benchmark instances are challenging, diverse, and rigorously validated.
@@
Shopping experts cohort, selection and input. We engaged two teams: 35 vetted domain experts contributing over 1,000 person-hours and a 15-member annotation team contributing over 3,000 person-hours. Experts curated domain-specific queries with chairs resolving disagreements, while annotators validated correctness, supplemented answers, and collected supporting evidence under expert guidance. This two-stage workflow integrates expert judgment with scalable verification.
3.3 Dataset Diversity and Human Time Cost
@@
ShoppingComp consists of 145 instances covering 558 scenarios, including 55 synthesized user instances, 42 expert-generated instances, and 48 safety critical instances. The benchmark follows the Amazon taxonomy and spans eleven categories with diverse task difficulties defined by constraint complexity, ranging from simple attribute-based queries to highly complex multi-constraint scenarios (see Appendix B).
@@
The time distribution for experts and annotators is shown in Figure 5. We collect response times from both experts and annotators through an online examination. To ensure fair comparison and reduce the impact of anomalous durations, we cap the recorded time: responses longer than one hour for experts and two hours for annotators are excluded from the final analysis. The long completion time suggests the intrinsic difficulty of ShoppingComp tasks, even for humans, establishing a rigorous upper bound for model evaluation.